Y varies directly as x and inversely as the square of z. y=24 when x= 12 and z=2. Find y when x=64 and z=4.
y=

Answer by richwmiller(17219)

y=kx
y=m/z
24=k*12 k=2
24=m/2 m=48
y=2*64=128 when x=64
y=48/4=12 when z=4